OpenStudy (anonymous):

Fan and Medal to best answer :) (-6)^-2?

OpenStudy (anonymous):

(-6)^-2 = 1/(-6)^2 = 1/36

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

yw